# Multimodal 3D Genome Pre-training

### v1

@@ 4.1 Self-supervised Pre-training
$\mathcal{L}_{\text{con}}=\frac{1}{2}(\mathcal{L}_{\text{pair}}(\hat{X_{E}^{I}},\hat{X_{M}^{I}})+\mathcal{L}_{\text{pair}}(\hat{X_{M}^{I}},\hat{X_{E}^{I}})).$ (2)
Moreover, we introduce orthogonal constraint to maximize the dissimilarity between modal-specific and their modal-invariant features. This ensures that the modal-specific features capture complementary information distinct from the shared knowledge represented by the modal-invariant features. The orthogonal loss $\mathcal{L}_{\text{orth}}$ is calculated by minimizing the inner product between these features as follows:
$\mathcal{L}_{\text{orth}}=\frac{1}{2}\left(\langle\hat{X_{M}^{S}},\hat{X_{M}^{I}}\rangle+\langle\hat{X_{E}^{S}},\hat{X_{E}^{I}}\rangle\right).$ (3)
Cross-modal mapping block. Data scarcity, often due to high experimental costs, can lead to incomplete datasets with missing modalities. Integrating predicted features of a missing modality with existing modalities can enhance prediction performance (Kim & Kim, 2025; Wang et al., 2023). Our cross-modal mapping block aims to capture the implicit semantic relationships and facilitate knowledge transfer between modalities to address this issue.
Some downstream tasks require the MIX-HIC to preserve the sequence length of the input features. However, the differing lengths of Hi-C contact maps and epigenomic tracks pose a challenge for effective modality transfer. Therefore, we apply 1D adaptive pooling to the concatenated representations of Hi-C, $X_{M}^{\text{Concat}}=[X_{M}^{I}:X_{M}^{S}]\in\mathbb{R}^{\alpha_{3}\times C_{3}}$ and epigenomic tracks, $X_{E}^{\text{Concat}}=[X_{E}^{I}:X_{E}^{S}]\in\mathbb{R}^{\beta_{3}\times C_{3}}$ (where [:][$\cdot$:$\cdot$][ ⋅ : ⋅ ] denotes concatenation), to align their lengths. This yields the complementary representations $X_{\text{M2E}}\in\mathbb{R}^{\beta_{3}\times C_{3}}$ and $X_{\text{E2M}}\in\mathbb{R}^{\alpha_{3}\times C_{3}}$ as follows:
$X_{\text{M2E}}=\mathcal{F}_{\text{M2E}}(\mathcal{G}_{\text{M2E}}(X_{M}^{\text{Concat}})),$ (4)
$X_{\text{E2M}}=\mathcal{F}_{\text{E2M}}(\mathcal{G}_{\text{E2M}}(X_{E}^{\text{Concat}})),$ (5)
where $\mathcal{G}_{\text{M2E}}$ and $\mathcal{G}_{\text{E2M}}$ are 1D adaptive pooling operations. $\mathcal{F}_{\text{M2E}}$ and $\mathcal{F}_{\text{E2M}}$ denote dense layers with the same output dimensions matching their inputs. To ensure these mapped embeddings capture the relevant information from the target modality, we use an $\mathcal{L}_{2}$ constraint for regularization:
@@
4.2 Task-specific Fine-tuning
As depicted in Figure 2, MIX-HIC is a versatile framework capable of processing various kinds of inputs. Ideally, MIX-HIC takes both the Hi-C contact map and the epigenomic feature profile as inputs, extracting their concatenated features $X_{M}^{\text{Concat}}$ and $X_{E}^{\text{Concat}}$ similar to the feature extraction block used in the self-supervised pre-training stage. However, in real-world scenarios, certain modality may be absent for various unforeseen reasons. Leveraging the powerful representation ability of pre-training to capture implicit connections between bimodal data, MIX-HIC incorporates a cross-modal mapping block to complement the features of the missing modality using the information from the available modality. For example, when the Hi-C contact map is missing, MIX-HIC can infer the missing modality features $X_{\text{E2M}}$ from the concatenated epigenomic embedding $X_{E}^{\text{Concat}}$ using Eq. 5. For the Hi-C contact map prediction task, even only the epigenomic tracks are available, the corresponding contact map features $X_{\text{E2M}}$ can still be utilized for prediction.
@@ 4.2 Task-specific Fine-tuning
Task-specific prediction block. Three types of decoders are involved in the task-specific prediction block. The chromatin loop prediction task is a binary classification problem, where the goal is to classify a given sample as either a chromatin loop or a non-loop. Therefore, the decoder for this task includes a mean pooling operation along the sequence length axis, followed by a feedforward network.
@@
Training loss. The chromatin loop prediction task employs the binary cross entropy (BCE) loss function, while the Hi-C contact map prediction and CAGE-seq expression prediction tasks use the mean squared error (MSE) loss function. These loss functions can be formulated as follows:
$\begin{gathered}\mathcal{L}_{\text{MSE}}=\frac{1}{J}\sum_{j=1}^{J}(y_{j}-\hat{y}_{j})^{2},\\ \mathcal{L}_{\text{BCE}}=-\frac{1}{J}\sum_{j=1}^{J}\left[y_{j}\log(\hat{y}_{j})+(1-y_{j})\log(1-\hat{y}_{j})\right].\end{gathered}$ (8)
@@
Table 2: Methods comparison for the Hi-C contact map prediction task on GM12878 and K562 cell lines. The results marked in bold and underlined denote the best and second-best performing methods, respectively.
@@
PCC $R^{2}$ PCC $R^{2}$
Epiphany 0.9159 0.7970 0.8629 0.6547
C.Origami 0.9258 0.7958 0.8934 0.7055
EPCOT-LSTM 0.9376 0.7993 0.9111 0.7840
EPCOT-transformer 0.9391 0.5409 0.9158 0.7648
MIX-HIC-InferMap 0.9441 0.8724 0.9181 0.8001
5 Experiments
Implementation details and hyperparameter analysis can be referred to Appendix.
5.1 Comparison Results on Downstream Tasks
To demonstrate the effectiveness of MIX-HIC, three downstream tasks are involved in this work, including Hi-C contact maps prediction, chromatin loops prediction, and CAGE-seq expression prediction.
Table 3: Methods comparison for the supervised chromatin loops prediction task on GM12878 and K562 cell lines.
@@ 5.1 Comparison Results on Downstream Tasks
Peakachu 0.7763 0.8283 0.8015 0.8766 0.7857 0.8062 0.7958 0.8864
DLoopCaller 0.8516 0.8130 0.8319 0.9118 0.8383 0.7526 0.7932 0.8924
MIX-HIC-Bimodal 0.8505 0.8337 0.8420 0.9209 0.8521 0.8027 0.8267 0.9194
3D Chromatin Organization Prediction. In this task, the Hi-C contact maps serve as the target. MIX-HIC-InferMap (a variant of MIX-HIC capable of inferring features of missing contact maps) is compared with four state-of-the-art methods, including Epiphany (Yang et al., 2023), C.Origami (Tan et al., 2023), and two variants of EPCOT (Zhang et al., 2023), (i.e. EPCOT-LSTM and EPCOT-transformer). For a fair comparison, all methods receive the same input as MIX-HIC and are configured with their default parameter settings. The coefficient of determination ($R^{2}$) (Nagelkerke, 1991) and the Pearson correlation coefficient (PCC) (Cohen et al., 2009) are employed as evaluation metrics.
Table 2 shows the evaluation performance on GM12878 and K562 cell lines. Compared to other methods, MIX-HIC demonstrates superior performance, achieving the highest average $R^{2}$ and PCC values on both GM12878 and K562 cell lines. Specifically, it outperforms the runner-up method by approximately 0.5%/9.1% in PCC/$R^{2}$ on GM12878, and by about 0.3%/2.1% on K562, respectively. We note that the EPCOT variants deliver competitive performance, but they suffer from fluctuations across the two datasets. Through extensive pre-training on large-scale pairwise datasets, MIX-HIC effectively explores implicit semantic relationships between bimodal data, enabling robust compensation for missing modality semantics and achieving promising accuracy in Hi-C contact map prediction.
Figure 3: Comparison of the number of predicted loops with the number of ChIA-PET-supported loops across various deep learning methods on GM12878 cell line.
Chromatin Loops Prediction. Two supervised machine learning-based methods DLoopCaller (Wang et al., 2022) and Peakachu (Salameh et al., 2020) are used to compare with MIX-HIC-Bimodal (a variant of MIX-HIC combining both Hi-C contact maps and epigenomic tracks). As illustrated in Table 3, MIX-HIC surpasses other machine learning-based methods across most classification metrics on two datasets. By integrating both epigenomic tracks and Hi-C contact maps, DLoopCaller slightly outperforms Peakachu, which relies solely on Hi-C contact maps. The enhanced performance of MIX-HIC can be attributed to the effective self-supervised pre-training, which facilitates the capture of both modal-invariant and modal-specific information from bimodal representations, in contrast to the simple concatenation employed by DLoopCaller.
We employ both statistical-based methods, namely ChromoSight (Matthey-Doret et al., 2020) and HiCExplorer (Wolff et al., 2020), as well as machine learning-based methods to annotate loops across entire chromosomes. The predicted loops are further compared with those loops validated by experimental data from ChIA-PET. Method details for whole-chromosome chromatin loop annotation and corresponding results are available in Appendix A.2 and C.1, respectively. Figure 3 compares the number of predicted loops versus the number of loops supported by ChIA-PET among various methods on GM12878 dataset. Among all the methods, machine learning-based methods generally predict a higher number of loops that are validated by ChIA-PET on the GM12878 dataset, with MIX-HIC identifying the most ChIA-PET validated loops while maintaining the highest validated proportion. On the K562 dataset, which contains much less training data compared to GM12878, the efficacy of machine learning methods generally declines. Nevertheless, MIX-HIC consistently maintains the highest proportion of validated loops in comparison with other methods. An example of whole-chromosome chromatin loop comparison is provided in Appendix C.2.
Table 4: Methods comparison for the CAGE-seq expression prediction task on GM12878 and K562 cell lines.
@@
PCC $R^{2}$ PCC $R^{2}$
EPI-CNN 0.9213 0.7719 0.9354 0.8033
EPI-Graph 0.9200 0.7965 0.9256 0.8211
EPCOT-LSTM 0.8939 0.4723 0.9359 0.8972
EPCOT-transformer 0.9337 0.8672 0.8965 0.7744
MIX-HIC 0.9435 0.8833 0.9569 0.9077
CAGE-seq Expression Prediction. MIX-HIC is pitted against four benchmark methods: two variants of GraphReg (EPI-CNN and EPI-Graph) (Karbalayghareh et al., 2022), as well as two variants of EPCOT (EPCOT-LSTM and EPCOT-transformer) (Zhang et al., 2023). EPI-CNN relies solely on epigenomic tracks to predict CAGE-seq expression, whereas EPI-Graph enhances this approach by employing graph attention networks to integrate features from Hi-C contact maps, following the feature extraction of epigenomic tracks. Table 4 showcases the comparison results across two datasets. EPCOT achieves competitive performance, yet it suffers from a time-consuming problem due to the process of long-range DNA sequence. MIX-HIC demonstrates remarkable performance over other benchmark methods across all metrics on both datasets.
@@ 5.2 Few-shot Chromatin Loops Classification
We conduct a few-shot learning experiment on the chromatin loops classification task to evaluate the performance of MIX-HIC under limited training data scenarios. Specifically, we fine-tune MIX-HIC, a non-pre-trained variant of MIX-HIC, and DLoopCaller, using four different ratios of the training data, ranging from 0.0001 to 0.1. As illustrated in Figure 4, MIX-HIC consistently outperforms the non-pre-trained model and DLoopCaller across all data ratios, demonstrating the significant advantage of pre-training. DLoopCaller exhibits considerable performance fluctuations when the training data is limited. Notably, with a training data ratio of 0.1, MIX-HIC achieves an AUROC of approximately 0.9 on two datasets, which is competitive with other state-of-the-art methods trained on full datasets. These findings highlight the robustness and efficiency of MIX-HIC in leveraging pre-trained knowledge to achieve superior performance even with limited labels.
Figure 5: Cross-cell-type evaluation of MIX-HIC performance on Hi-C contact map prediction and chromatin loops prediction tasks.

### v2

@@ 4.1 Self-supervised Pre-training
$).\mathcal{L}_{\text{con}}=\frac{1}{2}(\mathcal{L}_{\text{pair}}(\hat{X_{E}^{I}},\hat{X_{M}^{I}})+\mathcal{L}_{\text{pair}}(\hat{X_{M}^{I}},\hat{X_{E}^{I}}$)). (3)
Moreover, we introduce an orthogonal constraint to maximize the dissimilarity between modal-specific and modal-invariant features. This ensures that the modal-specific features capture complementary information distinct from the shared knowledge represented by the modal-invariant features. The orthogonal loss $\mathcal{L}_{\text{orth}}$ is calculated by minimizing the inner product between these features:
[…]
Cross-modal mapping block. Data scarcity, often due to high experimental costs, can lead to incomplete datasets with missing modalities. Integrating predicted features of a missing modality with existing modalities can enhance prediction performance wang2023multi ; kim2025missing . Our cross-modal mapping block aims to capture the implicit semantic relationships and facilitate knowledge transfer between modalities to address this issue.
Some downstream tasks require the MIX-HIC to preserve the sequence length of the input features. However, the differing lengths of Hi-C contact maps and epigenomic tracks pose a challenge for effective modality transfer. Therefore, we apply 1D adaptive pooling to the concatenated representations of Hi-C, $X_{M}^{\text{Concat}}=[X_{M}^{I}:X_{M}^{S}]\in\mathbb{R}^{\alpha_{3}\times C_{3}}$ and epigenomic tracks, $X_{E}^{\text{Concat}}=[X_{E}^{I}:X_{E}^{S}]\in\mathbb{R}^{\beta_{3}\times C_{3}}$ (where $\cdot$$\cdot$] denotes concatenation), to align their lengths. This yields the complementary representations $X_{\text{M2E}}\in\mathbb{R}^{\beta_{3}\times C_{3}}$ and $X_{\text{E2M}}\in\mathbb{R}^{\alpha_{3}\times C_{3}}$ as follows:
$X_{\text{M2E}}=\mathcal{F}_{\text{M2E}}(\mathcal{G}_{\text{M2E}}(X_{M}^{\text{Concat}})),X_{\text{E2M}}=\mathcal{F}_{\text{E2M}}(\mathcal{G}_{\text{E2M}}(X_{E}^{\text{Concat}})),$ (5)
where $\mathcal{G}_{\text{M2E}}$ and $\mathcal{G}_{\text{E2M}}$ are 1D adaptive pooling operations. $\mathcal{F}_{\text{M2E}}$ and $\mathcal{F}_{\text{E2M}}$ denote dense layers with the same output dimensions matching their inputs. To ensure these mapped embeddings capture the relevant information from the target modality, we use a $\mathcal{L}_{\text{mapping}}$ loss for regularization:
@@
4.2 Task-specific Fine-tuning
As depicted in Figure 2, MIX-HIC is a versatile framework capable of processing various kinds of inputs. Ideally, MIX-HIC takes both the Hi-C contact map and the epigenomic feature profile as inputs (represented as MIX-HIC-Bimodal), extracting their concatenated features $X_{M}^{\text{Concat}}$ and $X_{E}^{\text{Concat}}$ similar to the feature extraction block used in the self-supervised pre-training stage. However, in real-world scenarios, a certain modality may be absent for various unforeseen reasons.
@@ 4.2 Task-specific Fine-tuning
Leveraging the powerful representation ability of pre-training to capture implicit connections between bimodal data, MIX-HIC incorporates a cross-modal mapping block to complement the features of the missing modality using the information from the available modality. For example, when the Hi-C contact map is missing, MIX-HIC can infer the missing modality features $X_{\text{E2M}}$ from the concatenated epigenomic embedding $X_{E}^{\text{Concat}}$ using Eq. 5. For the Hi-C contact map prediction task, even if only the epigenomic tracks are available, the corresponding contact map features $X_{\text{E2M}}$ can still be utilized for prediction (denoted as MIX-HIC-Infer).
@@
Task-specific prediction block. Three types of decoders are involved in the task-specific prediction block. The chromatin loop detection task is a binary classification problem. Therefore, the decoder for this task includes a mean pooling operation along the sequence length axis, followed by a feedforward network, to classify a given sample as either a chromatin loop or a non-loop.
@@
Training loss. The chromatin loop detection task employs the binary cross entropy (BCE) loss function, while the Hi-C contact map prediction and CAGE-seq expression prediction tasks use the mean squared error (MSE) loss function. It should be noted that we normalize CAGE-seq data using RPGC ramirez2016deeptools2 and Hi-C data using KR kaul2020identifying normalization to correct for sequencing depth and systematic biases, respectively. Subsequently, a log transformation is applied to both datasets. This is critical for stabilizing variance and compressing the highly skewed distribution of raw counts, ultimately producing continuous values representing normalized interaction frequencies. Consequently, the MSE loss becomes an appropriate and robust choice for these continuous values. The MSE and BSE loss functions can be formulated as follows:
$\mathcal{L}_{\text{MSE}}=\frac{1}{J}\sum_{j=1}^{J}(y_{j}-\hat{y}_{j})^{2},$ (8)
$\mathcal{L}_{\text{BCE}}=-\frac{1}{J}\sum_{j=1}^{J}\left[y_{j}\log(\hat{y}_{j})+(1-y_{j})\log(1-\hat{y}_{j})\right],$ (9)
@@
5 Experiments
Implementation details, hyperparameter analysis, and analyses of the model’s biological grounding and robustness for noise can be referred to Appendix.
5.1 Comparison Results on Downstream Tasks
To demonstrate the effectiveness of MIX-HIC, three downstream tasks are involved in this work, including Hi-C contact map prediction, chromatin loop detection, and CAGE-seq expression prediction. We construct three versions of MIX-HIC: (1) MIX-HIC-Bimodal, which leverages both Hi-C contact maps and epigenomic tracks through pre-training; (2) MIX-HIC-NonPre, a non-pretrained version using the same bimodal inputs; and (3) MIX-HIC-Infer, designed to handle missing Hi-C data by integrating epigenomic track embeddings with inferred Hi-C embeddings. Detailed descriptions of the compared state-of-the-art methods are given in Appendix D.
Table 2: Methods comparison for the Hi-C contact map prediction task on GM12878 and K562 cell lines using $R^{2}$. The results marked in bold and underlined denote the best and second-best performing methods, respectively.
@@ 5.1 Comparison Results on Downstream Tasks
Epiphany yang2023epiphany 0.7970 0.6547
C.Origami tan2023cell 0.7958 0.7055
EPCOT-Transformer zhang2023generalizable 0.5409 0.7648
EPCOT-LSTM zhang2023generalizable 0.7993 0.7840
[…]
3D Chromatin Organization Prediction. In this task, the Hi-C contact maps serve as the target. MIX-HIC-Infer is compared with four state-of-the-art methods, including Epiphany yang2023epiphany , C.Origami tan2023cell , and two variants of EPCOT zhang2023generalizable , (i.e. EPCOT-LSTM and EPCOT-Transformer). For a fair comparison, all methods receive the same input as MIX-HIC and are configured with their default parameter settings. The coefficient of determination ($R^{2}$) nagelkerke1991note is employed as the evaluation metric, since it effectively quantifies explained variance in analyses of long-tailed or sparse data, such as low-probability long-range interactions in Hi-C.
Table 2 shows the evaluation performance on GM12878 and K562 cell lines. Compared to other methods, MIX-HIC-Infer demonstrates superior performance, achieving the highest average $R^{2}$ values on both GM12878 and K562 cell lines. Specifically, it outperforms the runner-up method by approximately 9.3% and 2.1% in $R^{2}$ score on GM12878 and K562, respectively. We note that the EPCOT variants deliver competitive performance, but they suffer from fluctuations across the two datasets. Through extensive pre-training on large-scale pairwise datasets, MIX-HIC effectively explores implicit semantic relationships between bimodal data, enabling robust compensation for missing modality semantics and achieving promising accuracy in Hi-C contact map prediction.
Chromatin Loop Detection. Two supervised machine learning-based methods DLoopCaller wang2022dloopcaller and Peakachu salameh2020supervised are used to compare with MIX-HIC-Bimodal. As illustrated in Table 3, MIX-HIC-Bimodal surpasses other machine learning-based methods across all classification metrics on two datasets. By integrating both epigenomic tracks and Hi-C contact maps, DLoopCaller slightly outperforms Peakachu, which relies solely on Hi-C contact maps. The enhanced performance of MIX-HIC can be attributed to the effective self-supervised pre-training, which facilitates the capture of both modal-invariant and modal-specific information from bimodal representations, in contrast to the simple concatenation employed by DLoopCaller.
Figure 3: Comparison of the number of predicted loops with the number of corresponding ChIA-PET-supported loops across various deep learning methods on GM12878 cell line.
We employ both statistical-based methods, namely ChromoSight matthey2020computer and HiCExplorer wolff2020galaxy , as well as machine learning-based methods to annotate loops across entire chromosomes. The predicted loops are further compared with those loops validated by experimental data from ChIA-PET. Method details for whole-chromosome chromatin loop annotation and corresponding results with the quantitative ‘proportion’ metric are available in Appendix A.2 and C.1, respectively. Figure 3 compares the number of predicted loops versus the number of loops supported by ChIA-PET among various methods on the GM12878 dataset. Among all the methods, machine learning-based methods generally predict a higher number of loops that are validated by ChIA-PET on the GM12878 dataset, with MIX-HIC identifying the most ChIA-PET validated loops while maintaining the highest validated proportion. On the K562 dataset, which contains much less training data compared to GM12878, the efficacy of machine learning methods generally declines. Nevertheless, MIX-HIC consistently maintains the highest proportion of validated loops in comparison with other methods. An example of whole-chromosome chromatin loop comparison is provided in Appendix C.2.
Table 3: Methods comparison for the supervised chromatin loop detection task on GM12878 and K562 cell lines.
@@
Peakachu salameh2020supervised 0.7763 0.8283 0.8015 0.8766 0.7895 0.7905 0.7900 0.8834
DLoopCaller wang2022dloopcaller 0.8433 0.8075 0.8250 0.9046 0.8383 0.7526 0.7932 0.8924
[…]
Table 4: Methods comparison for the CAGE-seq expression prediction task on GM12878 and K562 cell lines using $R^{2}$.
@@
EPI-CNN karbalayghareh2022chromatin 0.7719 0.8033
EPI-Graph karbalayghareh2022chromatin 0.7965 0.8211
EPCOT-LSTM zhang2023generalizable 0.4723 0.8704
EPCOT-Transformer zhang2023generalizable 0.8578 0.8230
[…]
CAGE-seq Expression Prediction. MIX-HIC-Bimodal is pitted against four benchmark methods: two variants of GraphReg (EPI-CNN and EPI-Graph) karbalayghareh2022chromatin , as well as two variants of EPCOT (EPCOT-LSTM and EPCOT-Transformer) zhang2023generalizable . EPI-CNN relies solely on epigenomic tracks to predict CAGE-seq expression, whereas EPI-Graph enhances this by incorporating Hi-C contact maps via graph attention networks for dual-modal modeling. Table 4 showcases the comparison results across two datasets. EPCOT achieves competitive performance, yet it suffers from a time-consuming problem due to the process of long-range DNA sequence. MIX-HIC demonstrates remarkable performance over other benchmark methods across all metrics on both datasets.
Overall, MIX-HIC outperforms state-of-the-art methods across all downstream tasks, with the most significant improvement (9.3% $R^{2}$) in Hi-C contact map prediction. Large-scale pre-training enables robust semantic representations for MIX-HIC, while other methods lack generalizability due to being narrowly optimized for specific tasks and often show inconsistent performance across datasets.
@@ 5.2 Few-shot Chromatin Loops Classification
We conduct a few-shot learning experiment on the chromatin loops classification task to evaluate the performance of MIX-HIC under limited training data scenarios. Specifically, we fine-tune MIX-HIC-Bimodal, MIX-HIC-NonPre, DLoopCaller, and Peakachu using four different ratios of the training data, ranging from 0.0001 to 0.1. As illustrated in Figure 4, MIX-HIC-Bimodal consistently outperforms other methods across all data ratios. Notably, the performance of most methods remains relatively robust even with minimal data, primarily stemming from the powerful and repetitive biological signatures of chromatin loops, which make the task tractable for most models. The performance variations among different architectures also highlight their inherent data efficiency. For instance, Peakachu, a random forest model, is inherently stable as it relies on engineered features that are less sensitive to data volume. In contrast, DLoopCaller’s CNN architecture is more data-hungry and thus shows a steeper decline. Our MIX-HIC architecture, even without pre-training, proves more data-efficient: the self-attention mechanism is better suited for capturing the global and long-range dependencies in contact maps than local CNNs, and its bimodal input provides complementary information, enhancing robustness even in low-data regimes.
With a training data ratio of 0.1, MIX-HIC-Bimodal achieves an AUROC of about 0.9 on two datasets, which is competitive with other state-of-the-art methods trained on full datasets. These findings highlight the robustness and efficiency of MIX-HIC in leveraging pre-trained knowledge to achieve superior performance even with limited labels.
Figure 5: Methods comparison for cross-cell-type evaluation.